OpenStudy (anonymous):

Find all the rational roots of the polynomial. x^3-3x^2-2x+6 I kind of know how to do this but I'm still confused about how to do this kind of problem.

OpenStudy (anonymous):

I know you use r/s

OpenStudy (anonymous):

and r=factors of constant and s=factors of leading coefficient

OpenStudy (anonymous):

(x^2-2)(x-3)

OpenStudy (anonymous):

x = 3 x = SQRT(2) x = -SQRT(2)

OpenStudy (cwrw238):

f(3) = 27 - 27 - 6 + 6 = 0 so (x-3) is a factor diiding by x-3 gives (x^2 - 2) so rots are 3 , sqrt2 and - sqrt2

OpenStudy (anonymous):

That is "trial and error". Better to look for factors of 6 such that you get -3x^2 and -2x.

OpenStudy (cwrw238):

i used factors of 6 : first i tried 2 than 3

OpenStudy (anonymous):

If you do, then you get (x^2-2)(x-3) = 0. That shows 3 is a factor.

OpenStudy (anonymous):

those arent the answers according to my hw site btw...

OpenStudy (anonymous):



OpenStudy (anonymous):

What are the answers according to your site?

OpenStudy (anonymous):

It doesn't give me the answers... I have 100 tries. lol

OpenStudy (anonymous):

It just said it's wrong

myininaya (myininaya):

\[x^2(x-3)-2(x-3)=0\] \[(x-3)(x^2-2)=0\] trail factors not needed

OpenStudy (anonymous):

Apparently, 3, sqrt(2) and -sqrt(2) are wrong answers. I can't figure out how they can be wrong.

OpenStudy (cwrw238):

trail factors?

OpenStudy (anonymous):

"trial" I think

myininaya (myininaya):

did you post the question right mario?

OpenStudy (anonymous):

yes

myininaya (myininaya):

and entering in: \[3, \sqrt{2}, -\sqrt{2}\] is not showing you are correct? Does it say anything else like round answers to nearest tenths or whatever?